Proposed FCAL Electronics R&D

Work on FCAL FE electronics will being summer 1993

• Use Central Calorimeter Digital Storage and Readout Card • Decide on packaging and necessary trigger signals

Resources required for next year are modest

• ∼ 1 man-month of design

• SDC portable DAQ for readout of Prototype HADCAL